Assignment Gradebook ORION Exercise D2-14 Complete each of the following contribution format income statements by supplying the missing numbers. b. d. Sales revenue $425,000 $700,000 Variable expenses 210,000 86,000 85,000 150,000 400,000 Fixed expenses Operating income Income taxes 70,000 120,000 18,000 16,00055,000 $10,500 $45,000 $145,00o LINK TO VIDEO TO TEXT Question Attempts: 0 of 3 used SAVE FORLATER SUBNIT ANS
Answer #1

a.

b.

c.

d.

Sales revenue

$    2,95,000.00

$ 4,25,000.00

$    2,67,000.00

$ 7,00,000.00

Variable expenses

$    2,10,000.00

$ 2,75,000.00

$        86,000.00

$ 3,00,000.00

Contribution margin

$        85,000.00

$ 1,50,000.00

$    1,81,000.00

$ 4,00,000.00

Fixed expenses

$        58,000.00

$      70,000.00

$    1,20,000.00

$ 2,00,000.00

Operating income

$        27,000.00

$      80,000.00

$        61,000.00

$ 2,00,000.00

Income taxes

$        16,500.00

$      18,000.00

$        16,000.00

$      55,000.00

Net income

$        10,500.00

$      62,000.00

$        45,000.00

$ 1,45,000.00

Working

a.

b.

c.

d.

Sales revenue

=210000+85000

425000

=181000+86000

700000

Variable expenses

210000

=425000-150000

86000

=700000-400000

Contribution margin

85000

150000

=120000+61000

400000

Fixed expenses

=85000-27000

70000

120000

=400000-200000

Operating income

27000

=150000-70000

=45000+16000

=145000+55000

Income taxes

=27000-10500

18000

16000

55000

Net income

10500

=80000-18000

45000

145000